A length of 20-gauge copper wire (of diameter 0.8118 mm) is formed into a circular loop with a radius of 23.0 cm. A magnetic field perpendicular to the plane of the loop increases from zero to 10.0 mT in 0.22 s. Find the average electrical power dissipated in the process.
Answer #1

resistance of wire

R = rho * L/ (pi/4 d^2)

R = 1.7* 10^-8* 2* 3.14* 0.23 / ( 3.14 / 4* (0.8118* 10^-3)^2)

R = 0.04746 ohm

now

induced emf

e = BA / t = 10* 10^-3* 3.14* 0.23^2 / 0.22

e = 7.55* 10^-3 V

power dissipation

P = e^2/R

P = 1.2* 10^-3 W

=======

Comment before rate in case any doubt, will reply for sure.. goodluck